# Pre-trained Universal Medical Image Transformer

### 12 Dec 2023 v1

@@ Results
We use the following two datasets to evaluate the segmentation performance. 1) The Beyond the Cranial Vault (BTCV) dataset, which contains 30 CT scans with 13 classes of regions annotated. We follow the train/validation split of SMIT (Jiang et al. 2022) to use 21 scans for training and 9 scans for validation. 2) The MRI subset of Combined Healthy Abdominal Organ Segmentation (CHAOS) dataset, which contains 40 pairs of T1-DUAL MRI images and 40 T2-SPIR MRI images with 4 classes of regions annotated. The segmentation labels of 20 pairs of T1-DUAL images and 20 T2-SPIR images are available for training, while the labels for remaining cases are hidden for online testing.
@@
For both tasks, we resample the images to the median spacing among the training data. Our pre-trained ViT is used to initialize the backbone, which extracts feature maps with the fixed in-plane scale of $\frac{1}{16}$. To build hierarchical feature maps for the segmentation head, following the implementation of EVA-02 (Fang et al. 2023a), we process the extracted feature maps with transposed convolutions to obtain feature maps with the in-plane scales of $\frac{1}{8}$ and $\frac{1}{4}$, and a max-pooling layer to obtain a feature map with the in-plane scale of $\frac{1}{32}$. Subsequently, a plain U-Net decoder processes the hierarchical feature maps to obtain the segmentation output. All segmentation models for both tasks are trained for 50K steps and the final weights are used for evaluation. Dice coefficient, average symmetric surface distance (ASSD), and Hausdorff distance (HD) are used to evaluate results for both tasks. The details on experimental settings and evaluation metrics are given in the appendix.
@@
Dice ASSD HD Dice ASSD HD
@@
Table 3: Segmentation results on the validation set of BTCV and hidden test set of CHAOS. “TFS” indicates the model with a plane ViT backbone trained from scratch. The values of Dice are shown in percentages.
Table 3 compares the segmentation performance of our model with other models on the BTCV validation set and CHAOS hidden test set. Our model outperforms all other models in terms of all 3 metrics for both tasks. Notably, although EVA-02-B is pre-trained on natural images, fine-tuning from it could already outperform Swin UNETR (Tang et al. 2022) and UniMiSS that are pre-trained on medical images. We hypothesize that this is due to the insufficient medical image data used for pre-training.
@@
Ablation Study
To investigate how our model improves the label efficiency of the downstream tasks, we compare the model trained from scratch and fine-tuned from our model with different ratios of training data in Table 4. Fine-tuning from our pre-trained model consistently outperforms training from scratch with a large margin across all metrics, and the improvements are more significant when less data are available. Moreover, our pre-trained model could effectively mitigate the requirement for labeled data. As shown in Table 4, even when fine-tuning with 60% of the training data, our pre-trained model still achieves better results than a model trained from scratch using the full dataset.
Conclusion
In this work, we introduce the SAC module, which serves as the building block for universal models for medical images. We are able to leverage 55 datasets for pre-training, which comprises over 9 million 2D slices (including over 48,000 3D images), representing the most extensive pre-training effort for 3D medical image models, to the best of our knowledge. Additionally, we develop a universal visual tokenizer for medical images, offering as a robust reconstruction objective for MIM. Our pre-trained model is demonstrated with superior downstream performance and label efficiency.
Appendix
Details on RoPE
From 1D to 2D RoPE
Positional embeddings play a crucial role in transformer models (Vaswani et al. 2017). RoPE (Su et al. 2021) is a kind of position embedding based on rotation transformation. We extend the 2D rotary position embedding (RoPE) (Su et al. 2021) to work with 3D medical images.
For a $d$-dimensional query or key vector $\bm{a}=(a_{0},a_{1},\dots,a_{d-1})$ in an attention block, the vanilla RoPE, primarily designed for sequence data, rotates each pair $(a_{2i},a_{2i+1})(i=0,1,\cdots,\frac{d}{2}-1)$ on the 2D plane by a radian of $\omega_{i}t$, where $t$ is the position index and it is commonly chosen that $\omega_{i}=b^{-\frac{2i}{d}}$ with a base $b$. This can be represented by matrix multiplication $\mathcal{R}(t)\bm{a}$, the transformation matrix is defined as
$\mathcal{R}(t)=\mathrm{diag}\left(T(\omega_{0}t),\dots,T(\omega_{\frac{d}{2}-1}t)\right),$ (7)
where $T(\theta)$ is defined as the 2D rotation matrix that rotates a radian of $\theta$:
$T(\theta)=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}.$ (8)
@@ Conclusion
  • Relativity. For two positions $(t_{x}^{(1)},t_{y}^{(1)})$ and $(t_{x}^{(2)},t_{y}^{(2)})$, it can be easily verified that $\mathcal{R}^{\mathsf{T}}(t_{x}^{(1)},t_{y}^{(1)})\mathcal{R}(t_{x}^{(2)},t_{y}^{(2)})=\mathcal{R}(t_{x}^{(2)}-t_{x}^{(1)},t_{y}^{(2)}-t_{y}^{(1)})$. Consequently, for two vectors $\bm{a}$ and $\bm{b}$, the result of their inner product after applying RoPE is a quantity dependent on the relative position $(t_{x}^{(2)}-t_{x}^{(1)},t_{y}^{(2)}-t_{y}^{(1)})$:
@@
  • Reversibility. $\mathcal{R}(t_{x}^{(1)},t_{y}^{(1)})=\mathcal{R}(t_{x}^{(2)},t_{y}^{(2)})$ if and only if $(t_{x}^{(1)},t_{y}^{(1)})=(t_{x}^{(2)},t_{y}^{(2)})$, as long as the value of each $\omega_{i}$ are appropriately chosen.
Additive 3D RoPE
@@
Relativity. We only need to show $\mathcal{R}_{x,z}(\cdot,\cdot)$ satisfies relativity. For each diagonal entry, we have:
$T^{\mathsf{T}}(\omega_{x,i}t_{x}^{(1)}+\omega_{z,i}t_{z}^{(1)})T(\omega_{x,i}t_{x}^{(2)}+\omega_{z,i}t_{z}^{(2)})$ (15)
$=$ $T(\omega_{x,i}(t_{x}^{(2)}-t_{x}^{(1)})+\omega_{z,i}(t_{z}^{(2)}-t_{z}^{(1)})).$ (16)
@@
Reversibility. Similarly, we only need to show $\mathcal{R}_{x,z}(\cdot,\cdot)$ satisfies reversibility. Note that the values we specifically choose for $\omega_{x,i},\omega_{y,i},\omega_{z,i}$ are algebraic numbers, the following theorem helps proves the result.
Theorem 0.1.
If both $\omega_{x}$ and $\omega_{z}$ are non-zero algebraic numbers, and $\frac{\omega_{x}}{\omega_{z}}$ is an irrational number, then for any $(t_{x}^{(1)},t_{z}^{(1)}),(t_{x}^{(2)},t_{z}^{(2)})\in\mathbb{Z}^{2}$ where $(t_{x}^{(1)},t_{z}^{(1)})\neq(t_{x}^{(2)},t_{z}^{(2)})$, it holds that $\mathcal{R}_{x,z}(t_{x}^{(1)},t_{z}^{(1)})\neq\mathcal{R}_{x,z}(t_{x}^{(2)},t_{z}^{(2)})$
@@
If $k=0$, then $t_{x}^{(2)}\neq t_{x}^{(1)}$ (otherwise, from $0=\omega_{x}(t_{x}^{(2)}-t_{x}^{(1)})+\omega_{z}(t_{z}^{(2)}-t_{z}^{(1)})$, we would also have $t_{z}^{(2)}=t_{z}^{(1)}$, which contradicts $(t_{x}^{(1)},t_{z}^{(1)})\neq(t_{x}^{(2)},t_{z}^{(2)})$). Therefore, from (18) we have:
@@
If $k\neq 0$, since the product and sum of algebraic numbers are still algebraic, from (18), we have
@@
Experiment
To validate the effectiveness of the proposed Additive 3D RoPE, we computed the rotation radians corresponding to all positions on images of a typical size. Note that the $x$ and $y$ dimensions are symmetric in Additive 3D RoPE, therefore, in our experiments, we set the length of the $y$ dimension to 1, and only focus on the $x$ and $z$ dimensions. Accordingly, for each $10\leq i\leq d/4-$ 1 where $d$ is the dimension of the query or key vector $\bm{a}$), we examined the distribution of these radians in terms of $\bmod 2\pi$ representation.
For image size to be validated, we set it as $8\times 16\times 1(z\times x\times y)$. Notice that that when the z-coordinate $t_{z}=0$, the rotation radian becomes $\omega_{x,i}t_{x}+\omega_{z,i}t_{z}=\omega_{x,i}t_{x}$, which is equivalent to 2D RoPE. Therefore, we also choose 2D image size of $1\times 16\times 1(z\times x\times y)$ for comparison with 2D RoPE.
For each $10\leq i\leq d/4-$ 1, we consider the results of rotating radians of all positions modulo $2\pi$ to obtain $\Theta_{i}$. Subsequently, we calculate the differences between all adjacent elements after sorting $\Theta_{i}$, resulting in $\Delta_{i}$. Finally, we calculate the minimum value among these differences as $\min\Delta_{i}$. The formal definitions of $\Theta_{i},\Delta_{i}$ are as follows:
@@
In addition, we also attempt to use the Euclidean distance between the positions after rotating $(0,1)\in\mathbb{R}^{2}$ by radians of $\alpha$ and $\beta$ as an indicator to quantify the ”difference” between these two rotation radians, as follows:
@@
Since there is a considerable number of $\theta_{i}$ values, the distribution of them on the unit circle $x^{2}+y^{2}=1$ is relatively dense. In this scenario, $\sqrt{(\cos\theta_{i+1}-\cos\theta_{i})^{2}+(\sin\theta_{i+1}-\sin\theta_{i})^{2}}\approx(\theta_{i+1}-\theta_{i})\bmod 2\pi$, implying that the arc length corresponding to this central angle will be approximately equal to the chord length. Consequently, the results obtained from these two evaluation metrics will be quite close. For simplicity, only the differences of adjacent radians are presented below.
Figure 4: Minimum intervals between rotation radians for different $i$. $i$ is the index of dimensions $d$ of query or key vector $\bm{a}$.
It can be observed that when using Additive 3D RoPE on 3D images, the minimum intervals between rotation radians are approximately in the range of $10^{-3}$ to $10^{-5}$, which preserves good numerical precision with 32bit floating point numbers. As for the comparison with 2D RoPE, the average result of $\frac{\min\Delta_{i}^{(2D)}}{\min\Delta_{i}^{(3D)}}$ for all $i$ is 26.30, which benefits 3D medical images processing. By setting $b_{z}<b_{x}$, which implies $\omega_{z,i}>\omega_{x,i}$, Additive 3D RoPE gains anisotropy, allowing it to carry cross-plane positional information. The effect of this choice on the distribution of rotation radians is that rotation radians for positions with the same $z$ coordinate tend to cluster together, while positions with different $z$ coordinates in the entire image give rise to distinct clusters of rotation radians. Thus, the difference between our results and the ideal ratio of 16.00 for both cases with uniform distribution reflects the anisotropic information.
As shown in Figure 4, we can also observe the relationship between $b_{z}$ and $\min\Delta_{i}$: if $b_{z}$ is large, indicating that $\omega_{z,i}$ is small, all rotation radians are densely distributed, resulting in a smaller $\min\Delta_{i}$; whereas if $b_{z}$ is small, although $\min\Delta_{i}$ is better, larger $\omega_{z,i}$ can lead to rotation radians for positions that are far apart become too close in terms $\bmod 2\pi$, which contradicts the relativity pursued by RoPE. In conclusion, we finally set $b_{z}=2333$ in our work.
@@
optimizer AdamW (Loshchilov and Hutter 2019)
@@
Pre-training Data and Pre-processing

### 24 Jan 2024 v2

@@ 5.2 Classification Experiments
MedMNIST v2 Yang et al. (2023) is a large-scale benchmark for medical image classification on standardized MNIST-like 2D and 3D images with diverse modalities, dataset scales, and tasks. The size of each image is either $28\times 28$ (2D) or $28\times 28\times 28$ (3D). We primarily experiment on the 3D portion of MedMNIST v2, namely the Organ Bilic et al. (2019); Xu et al. (2019), Nodule Armato III et al. (2011), Fracture Jin et al. (2020a), Adrenal, Vessel Yang et al. (2020), and Synapse datasets. We also experiment on 3 tasks from the 2D portion, namely the Derma Tschandl et al. (2018); Codella et al. (2019), Retina Liu et al. (2022), and Tissue Ljosa et al. (2012) datasets.
@@
Like common ViT-based models, our model performs classification tasks by feeding the hidden state of the [CLS] token into a fully-connected layer. We conduct classification experiments by fine-tuning from our and several other self-supervised pre-trained models on each dataset. Following the official split of training/validation/test sets, we train all models on the training sets for 100 epochs. The model weights from the epoch with the highest AUC score on the validation set are selected for comparison on the test set. The experimental details are given in the appendix.
Results
Table 2 compares the classification performance of our model with other models on the 3D portion using area under the ROC curve (AUC) and accuracy (ACC) metrics. Our model consistently outperforms all other models across all datasets in terms of AUC, and achieves the highest ACC on 4 out of 6 datasets. On average, our model improves AUC by at least 4.4% and accuracy by at least 3.4% across all datasets. Notably, on the Synapse dataset, our model improves the performance by a significantly large margin with 10.6% of AUC and 9.7% of ACC. The imaging modality of the Synapse dataset, electron microscope, is not covered by our pre-training data, showing the generalizability of our model. Table 3 shows the classification performance on 2D tasks. Our model outperforms all other baselines except for EVA-02 with a large margin, and is comparable to EVA-02. We hypothesize that the large-scale pre-training in general vision domain and the natural language supervision of EVA-02 can benefit 2D medical image classification; however, the results are still inferior on 3D images due to significant distribution shift.
Method Derma Retina Tissue Average
AUC ACC AUC ACC AUC ACC AUC ACC
ResNet-50 91.2 73.1 71.6 51.1 93.2 68.0 85.3 64.1
AutoML Vision 91.4 76.8 75.0 53.1 92.4 67.3 86.3 65.7
UniMiSS 90.9 74.1 71.2 52.5 92.8 67.6 85.0 64.7
EVA-02-B 94.5 80.1 77.8 54.5 95.1 73.5 89.1 69.4
Ours 94.6 79.2 77.4 55.3 94.9 73.0 89.0 69.2
Table 3: Classification results on the test sets of the 2D image datasets from MedMNIST v2. The results of ResNet-50 and Google AutoML Vision are taken from the original paper Yang et al. (2023). Both AUC and ACC values are shown in percentages.
5.3 Segmentation Experiments
Dataset
We use the following two datasets to evaluate the segmentation performance. 1) The Beyond the Cranial Vault (BTCV) dataset, which contains 30 CT scans with 13 classes of regions annotated. We follow the train/validation split of SMIT Jiang et al. (2022) to use 21 scans for training and 9 scans for validation. 2) The MRI subset of Combined Healthy Abdominal Organ Segmentation (CHAOS) dataset, which contains 40 pairs of T1-DUAL MRI images and 40 T2-SPIR MRI images with 4 classes of regions annotated. The segmentation labels of 20 pairs of T1-DUAL images and 20 T2-SPIR images are available for training, while the labels for remaining cases are hidden for online testing.
Implementation Details
For both tasks, we resample the images to the median spacing among the training data. Our pre-trained SPAD-ViT is used to initialize the backbone, which extracts feature maps with the fixed in-plane scale of $\frac{1}{16}$. To build hierarchical feature maps for the segmentation head, following the implementation of EVA-02 Fang et al. (2023a), we process the extracted feature maps with transposed convolutions to obtain feature maps with the in-plane scales of $\frac{1}{8}$ and $\frac{1}{4}$, and a max-pooling layer to obtain a feature map with the in-plane scale of $\frac{1}{32}$. Subsequently, a plain U-Net decoder processes the hierarchical feature maps to obtain the segmentation output. All segmentation models for both tasks are trained for 50K steps and the final weights are used for evaluation. Dice coefficient, average symmetric surface distance (ASSD), and Hausdorff distance (HD) are used to evaluate results for both tasks. The details on experimental settings and evaluation metrics are given in the appendix.
@@ Results
Dice ASSD $\downarrow$ HD $\downarrow$ Dice ASSD HD
@@
Table 4: Segmentation results on the validation set of BTCV and hidden test set of CHAOS. “TFS” indicates training a plane ViT backbone from scratch. The Dice values are shown in percentages.
Table 4 compares the segmentation performance of our model with other models on the BTCV validation set and CHAOS hidden test set. The models for comparison include Swin UNETR, the previous SOTA on MSD Antonelli et al. (2022). Our model consistently outperforms all other models in terms of all 3 metrics for both tasks. On the BTCV task, our model improves the HD metric by a large margin of $3.19$. On the CHAOS tasks, our model could still achieve evident improvement although the baseline performance is almost saturate.
@@
Few-shot Learning
To investigate how our model improves the label efficiency of the downstream tasks, we compare the model trained from scratch and fine-tuned from our model with different ratios of training data in Table 5. Fine-tuning from our pre-trained model consistently outperforms training from scratch with a large margin across all metrics, and the improvements are more significant when fewer data are available. Moreover, our pre-trained model could effectively mitigate the requirement for labeled data. As shown in Table 5, even when fine-tuning with 60% of the training data, our pre-trained model still achieves better results than a model trained from scratch using the full dataset. Notably, our model still preserves decent performance with only 4 labeled samples (20% of the training data).
6 Conclusion
In this work, we develop the first universal foundation model for medical image analysis, to our knowledge. To address the heterogeneous spatial properties of medical images, we introduce the concept of SPAD-Nets, a family of networks which dynamically adjust the structures to adapt to the spatial properties of input images. We build SPAD-ViT which is able to leverage 55 diverse datasets for pre-training, representing the most extensive pre-training effort for universal foundation models for medical image analysis to our knowledge. In addition, we develop SPAD-VT, a universal visual tokenizer for a broad range of medical images, offering a robust reconstruction objective for MIM, and may also be used in multimodal models in future works. The pre-trained SPAD-ViT is demonstrated with superior downstream performance and label efficiency. We hope our work could inspire future works on developing foundation models for medical image analysis.
Appendix A Details on RoPE
A.1 From 1D to 2D RoPE
Positional embeddings play a crucial role in transformer models Vaswani et al. [2017]. RoPE Su et al. [2021] is a kind of position embedding based on rotation transformation. We extend the 2D rotary position embedding (RoPE) Su et al. [2021] to work with 3D medical images.
For a $d$-dimensional query or key vector $\bm{a}=(a_{0},a_{1},\dots,a_{d-1})$ in an attention block, the vanilla RoPE, primarily designed for sequence data, rotates each pair $(a_{2i},a_{2i+1})$ (where $i\in\{0,1,\cdots,\frac{d}{2}-1\}$) on the 2D plane by an angle of $\omega_{i}t$ (in radian), where $t$ is the position index and it is commonly chosen that $\omega_{i}=b^{-\frac{2i}{d}}$ with a base $b$. This can be represented by matrix multiplication $\mathcal{R}(t)\bm{a}$, the transformation matrix is defined as
$\mathcal{R}(t)=\mathrm{diag}\left(R(\omega_{0}t),\dots,R(\omega_{\frac{d}{2}-1}t)\right),$ (4)
where $R(\theta)$ is defined as the 2D rotation matrix that rotates an angle of $\theta$:
$R(\theta)=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}.$ (5)
@@ A.1 From 1D to 2D RoPE
  1. 1.
    Relativity: for two positions $(t_{x}^{(1)},t_{y}^{(1)})$ and $(t_{x}^{(2)},t_{y}^{(2)})$, it can be easily verified that $\mathcal{R}^{\mathsf{T}}(t_{x}^{(1)},t_{y}^{(1)})\mathcal{R}(t_{x}^{(2)},t_{y}^{(2)})=\mathcal{R}(t_{x}^{(2)}-t_{x}^{(1)},t_{y}^{(2)}-t_{y}^{(1)})$. Consequently, for two vectors $\bm{a}$ and $\bm{b}$, the result of their inner product after applying RoPE is a quantity dependent on the relative position $(t_{x}^{(2)}-t_{x}^{(1)},t_{y}^{(2)}-t_{y}^{(1)})$:
@@
  2. 2.
    Reversibility: $\mathcal{R}(t_{x}^{(1)},t_{y}^{(1)})=\mathcal{R}(t_{x}^{(2)},t_{y}^{(2)})$ if and only if $(t_{x}^{(1)},t_{y}^{(1)})=(t_{x}^{(2)},t_{y}^{(2)})$, as long as the value of each $\omega_{i}$ are appropriately chosen.
A.2 Additive 3D RoPE
@@
Relativity
Obviously, we only need to show $\mathcal{R}_{x,z}(\cdot,\cdot)$ satisfies relativity. For each diagonal entry, we have:
$R^{\mathsf{T}}(\omega_{x,i}t_{x}^{(1)}+\omega_{z,i}t_{z}^{(1)})R(\omega_{x,i}t_{x}^{(2)}+\omega_{z,i}t_{z}^{(2)})$ (12)
$=$ $R(\omega_{x,i}(t_{x}^{(2)}-t_{x}^{(1)})+\omega_{z,i}(t_{z}^{(2)}-t_{z}^{(1)})).$ (13)
@@
Reversibility
Similarly, we only need to show $\mathcal{R}_{x,z}(\cdot,\cdot)$ satisfies reversibility. Note that the values we specifically choose for $\omega_{x,i},\omega_{y,i},\omega_{z,i}$ are algebraic numbers, the following theorem helps prove the result.
Theorem 1.
If both $\omega_{x}$ and $\omega_{z}$ are non-zero algebraic numbers, and $\frac{\omega_{x}}{\omega_{z}}$ is an irrational number, then for any $(t_{x}^{(1)},t_{z}^{(1)}),(t_{x}^{(2)},t_{z}^{(2)})\in\mathbb{Z}^{2}$ where $(t_{x}^{(1)},t_{z}^{(1)})\neq(t_{x}^{(2)},t_{z}^{(2)})$, it holds that $\mathcal{R}_{x,z}(t_{x}^{(1)},t_{z}^{(1)})\neq\mathcal{R}_{x,z}(t_{x}^{(2)},t_{z}^{(2)})$.
@@
If $k=0$, then $t_{x}^{(2)}\neq t_{x}^{(1)}$ (otherwise, from $0=\omega_{x}(t_{x}^{(2)}-t_{x}^{(1)})+\omega_{z}(t_{z}^{(2)}-t_{z}^{(1)})$, we would also have $t_{z}^{(2)}=t_{z}^{(1)}$, which contradicts $(t_{x}^{(1)},t_{z}^{(1)})\neq(t_{x}^{(2)},t_{z}^{(2)})$). Therefore, from Equation 15 we have:
@@
If $k\neq 0$, since the product and sum of algebraic numbers are still algebraic, from Equation 15, we have
@@
A.3 Experiment
To validate the effectiveness of the proposed Additive 3D RoPE, we computed the rotation angles corresponding to all positions on images of a typical size. Note that the $x$ and $y$ dimensions are symmetric in Additive 3D RoPE, therefore, in our experiments, we set the length of the $y$ dimension to 1, and only focus on the $x$ and $z$ dimensions. Accordingly, for each $10\leq i\leq d/4-$ 1 where $d$ is the dimension of the query or key vector $\bm{a}$), we examined the distribution of these angles in terms of $\bmod 2\pi$ representation.
For image size to be validated, we set it as $8\times 16\times 1(z\times x\times y)$. Notice that that when the z-coordinate $t_{z}=0$, the rotation angle becomes $\omega_{x,i}t_{x}+\omega_{z,i}t_{z}=\omega_{x,i}t_{x}$, which is equivalent to 2D RoPE. Therefore, we also choose 2D image size of $1\times 16\times 1(z\times x\times y)$ for comparison with 2D RoPE.
For each $10\leq i\leq d/4-$ 1, we consider the results of rotating angles of all positions modulo $2\pi$ to obtain $\Theta_{i}$. Subsequently, we calculate the differences between all adjacent elements after sorting $\Theta_{i}$, resulting in $\Delta_{i}$. Finally, we calculate the minimum value among these differences as $\min\Delta_{i}$. The formal definitions of $\Theta_{i},\Delta_{i}$ are as follows:
@@ A.3 Experiment
In addition, we also attempt to use the Euclidean distance between the positions after rotating $(0,1)\in\mathbb{R}^{2}$ by angles of $\alpha$ and $\beta$ as an indicator to quantify the ”difference” between these two rotation angles, as follows:
@@
Since there is a considerable number of $\theta_{i}$ values, the distribution of them on the unit circle $x^{2}+y^{2}=1$ is relatively dense. In this scenario, $\sqrt{(\cos\theta_{i+1}-\cos\theta_{i})^{2}+(\sin\theta_{i+1}-\sin\theta_{i})^{2}}\approx(\theta_{i+1}-\theta_{i})\bmod 2\pi$, implying that the arc length corresponding to this central angle will be approximately equal to the chord length. Consequently, the results obtained from these two evaluation metrics will be quite close. For simplicity, only the differences of adjacent angles are presented below.
Figure 5: Minimum intervals between rotation angles for different $i$. $i$ is the index of dimensions $d$ of query or key vector $\bm{a}$.
It can be observed that when using Additive 3D RoPE on 3D images, the minimum intervals between rotation angles are approximately in the range of $10^{-3}$ to $10^{-5}$, which preserves good numerical precision with 32bit floating point numbers. As for the comparison with 2D RoPE, the average result of $\frac{\min\Delta_{i}^{(2D)}}{\min\Delta_{i}^{(3D)}}$ for all $i$ is 26.30, which benefits 3D medical images processing. By setting $b_{z}<b_{x}$, which implies $\omega_{z,i}>\omega_{x,i}$, Additive 3D RoPE gains anisotropy, allowing it to carry cross-plane positional information. The effect of this choice on the distribution of rotation angles is that rotation angles for positions with the same $z$ coordinate tend to cluster together, while positions with different $z$ coordinates in the entire image give rise to distinct clusters of rotation angles. Thus, the difference between our results and the ideal ratio of 16.00 for both cases with uniform distribution reflects the anisotropic information.
As shown in Figure 5, we can also observe the relationship between $b_{z}$ and $\min\Delta_{i}$: if $b_{z}$ is large, indicating that $\omega_{z,i}$ is small, all rotation angles are densely distributed, resulting in a smaller $\min\Delta_{i}$; whereas if $b_{z}$ is small, although $\min\Delta_{i}$ is better, larger $\omega_{z,i}$ can lead to rotation angles for positions that are far apart become too close in terms $\bmod 2\pi$, which contradicts the relativity pursued by RoPE. In conclusion, we finally set $b_{z}=2333$ in our work.
@@
optimizer AdamW Loshchilov and Hutter [2019]
@@
Appendix B Interpretation of PDR
In fact, there exists a constructive interpretation when using PDR with the soft token representation. Suppose the encoder outputs a $D\times H\times W$ grid of token distributions, let $G$ denote the set of discrete coordinates of all cells within the grid. For each position $s\in G$ on the grid, we define a random variable $I_{s}$ denoting the index of the token in the codebook at $s$ and its distribution to be $q^{(s)}$. We define another random variable $S$ representing a random position on the grid that is uniformly distributed over $G$. It can be shown that the averaged distribution of all token distributions on the grid is exactly the distribution of $I_{S}$:
$=\frac{1}{DHW}\sum_{s\in G_{D,H,W}}\mathrm{P}(I_{s}=i).$ (18)
Intuitively, when the distribution of $I_{S}$ is close to the uniform distribution, individual tokens occur at a randomly sampled position with close probabilities, which benefits the diversity of the learned distributions.
Appendix C Pre-training Data and Pre-processing